OpenStudy (anonymous):

A negative charge of -2.0 x 10^-4C and a positive charge of 8.0 x 10^-4C are separated by 0.30m. What is the force between the two charges? please can someone help me with this problem

OpenStudy (anonymous):

you know coulomb's law right?

OpenStudy (anonymous):

yea kind of , not alot

OpenStudy (anonymous):

i dont know how to solve this problem tho

OpenStudy (anonymous):

but you do know the equation\[\large F=k\frac{q_1q_2}{r^{2}}\]right?

OpenStudy (anonymous):

yea i do

OpenStudy (anonymous):

so just put q1=2.0 x 10^-4C q2=8.0 x 10^-4C r=0.3m & find F. Note that i've ignored the sign of q1.

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

so then how do i find f?

OpenStudy (anonymous):

@rajathsbhat so what is the answer ?

OpenStudy (ujjwal):

@rajathsbhat has already answered it.. He gave you the formula and the value of variables used in formula.. Just substitute them in formula! Use K=9*10^9

OpenStudy (anonymous):

okay thank you :)

OpenStudy (anonymous):

so the answer wod be 6.6177 ???

OpenStudy (anonymous):

wait wat would K be ?

OpenStudy (ujjwal):

what formula are you using?

OpenStudy (anonymous):

the F=k*q1*q2 /r2

OpenStudy (anonymous):

bt i dont know wat k is ?

OpenStudy (shane_b):

This should help: http://hyperphysics.phy-astr.gsu.edu/hbase/electric/elefor.html

OpenStudy (anonymous):

thank you so much :)

OpenStudy (shane_b):

np :)